# Dissipative dynamics of an open quantum battery in the BTZ spacetime

### v1

@@ Abstract
We consider how charging performances of a quantum battery, modeled as a two-level system, are influenced by the presence of vacuum fluctuations of a quantum field satisfying the Dirichlet, transparent, and Neumann boundary conditions in the BTZ spacetime. The quantum battery is subjected to an external static driving which works as a charger. Meanwhile, the quantum field is assumed to be coupled to both longitudinal and transverse spin components of the quantum battery including decoherence and pure dephasing mechanisms. Charging and discharging dynamics of the quantum battery are derived by extending the previous open quantum system approach in the relativistic framework to this more general scenario including both the driving and multiple coupling. Analytic expressions for the time evolution of the energy stored are presented. We find that when the driving amplitude is stronger/weaker than the energy-level spacing of the quantum battery the pure dephasing dissipative coupling results in better/worse charging performances than the decoherence dissipative coupling case. We also find that higher Hawking temperature helps to improve the charging performance under certain conditions compared with the closed quantum buttery case, implying the feasibility of energy extraction from vacuum fluctuations in curved spacetime via dissipation in charging protocol. Different boundary conditions for quantum field may lead to different charging performance. Furthermore, we also address the charging stability by monitoring the energy behaviour after the charging protocol has been switched off. Our study presents a general framework to investigate relaxation effects in curved spacetime, and reveals how spacetime properties and field boundary condition affect the charging process, which in turn may shed light on the exploration of the spacetime properties and thermodynamics via the charging protocol.
@@ 1 Introduction
The notion of QB was first introduced in the seminal work PhysRevE.87.042123 by Alicki and Fannes, showing entangling unitary controls extract in general more work than independent ones. Afterward, the development of different figures of merit, such as the charging time, the associated power, and the amount of extractable work, has been a major focus on the QB physics Campaioli:2023ndh ; Binder_2015 ; PhysRevLett.122.047702 ; PhysRevLett.111.240401 ; PhysRevLett.111.050601 ; PhysRevLett.131.030402 ; PhysRevLett.129.130602 . To improve the charging performance, a lot of methods have been developed to assist the charging protocol, including collective quantum resources PhysRevLett.120.117702 ; PhysRevLett.118.150601 ; PhysRevLett.128.140501 , ordered and disordered interactions PhysRevA.101.032115 , external environment engineering PhysRevA.107.032203 ; PhysRevE.104.064143 , collisional model PhysRevLett.127.100601 ; PhysRevResearch.5.013155 , Sachdev-Ye-Kitaev model PhysRevLett.125.236402 ; SYK and so on; see Ref. Campaioli:2023ndh for an extensive review and a comprehensive list of references. Beyond the theoretical analysis, the experimental implementation of various kinds of QBs and different charging protocols has also been reported recently Campaioli:2023ndh ; PhysRevB.108.L180301 ; PhysRevLett.131.240401 ; Hu_2022 ; PhysRevA.106.042601 ; batteries8050043 . Furthermore, since any systems are unavoidably coupled to the environment, at least subjected to vacuum fluctuations at the quantum scale, relaxation and dephasing would play a key role in the system dynamics. In this sense, the QB’s figures of merit, including the maximum average energy stored during the charging protocol, the charging power, the charging stability and so on, inevitably undergoes the relaxation and dephasing, which may even dominate the QB charging performance. It is thus of great importance to learn the charging performance in the presence of inevitable external environmental action Campaioli:2023ndh ; PhysRevB.99.035421 ; PhysRevA.102.060201 ; PhysRevA.103.033715 ; Carrega_2020 ; PhysRevLett.122.210601 ; PhysRevE.104.044116 ; PhysRevA.104.032207 .
@@
The outline of this paper is as follows. In section 2 we present the model of a single cell QB coupled to a vacuum fluctuations of quantum field. The charging performance indicators, such as the average energy stored in the QB, are calculated here. In section 3 we consider the dynamics of the average energy stored when the open QB is located in the BTZ spacetime, and explore how the driving amplitude, BTZ spacetime properties, the boundary conditions on quantum field, and different dissipative couplings affect the charging performance, including both the dynamics of the average energy stored and the charging stability. Finally, discussions and summary of the main results are present in section 4.
@@ 2.1 Open quantum battery model
where $\mu$ is a small coupling constant. This interaction captures both decoherence ($\theta=0$) and pure dephasing ($\theta=\pi/2$) processes.
@@
Rotating $\tilde{\rho}_{\text{BQ}}(\tau)$ in Eq. (12) with state parameters (2.1), we can obtain the evolving state of the QB, $\rho_{\text{QB}}(\tau)=R^{\dagger}(\Theta)\tilde{\rho}_{\text{QB}}(\tau)R(\Theta)$.
@@ 2.2 Charging performance indicators
Note that $r_{3}(\tau)$ and $r_{1}(\tau)$ depend on the field correlation function shown in (2.1), thus on the spacetime properties due to the scattering off quantum field. Therefore, one can expect that the charging dynamics depends on the spacetime properties, and in turn one can also learn the spacetime properties from the charging physics.
@@ 3 Quantum battery in BTZ spacetime
To obtain the charging dynamics of the QB in the BTZ spacetime, we first need to know the Wightman function of the conformal scalar field which may capture the character of the quantum field vacuum fluctuations and plays an important role in the dynamics of the QB through the QB-field interaction. In this paper, we assume that the scalar field is in the Hartle-Hawking vacuum in the BTZ spacetime. It is interesting to note that the BTZ spacetime can be obtained by a topological identification of $\mathrm{AdS}_{3}$ spacetime. Therefore, using the method of images, the corresponding Wightman function in the BTZ spacetime for the conformal massless scalar field in the Hartle-Hawking vacuum can be obtained in terms of the corresponding Wightman function in $\mathrm{AdS}_{3}$ spacetime PhysRevD.49.1929 ,
@@
with $\Delta\phi=\phi-\phi^{\prime}$, and $\Delta\,t=t-t^{\prime}$. The parameter $\zeta\in\{1,0,-1\}$ respectively specifies the Dirichlet ($\zeta=1$), transparent ($\zeta=0$), and Neumann ($\zeta=-1$) boundary conditions satisfied by the field at spatial infinity PhysRevD.18.3565 .
@@
Let us note that the local temperature of the BTZ spacetime can be rewritten in the following form
@@
with $a_{\text{BTZ}}=r/(\ell\sqrt{r^{2}-r^{2}_{\text{h}}})$ representing the acceleration of the constant $r$ trajectory in the BTZ spacetime. Here $T$ is analogous to the temperature felt by an accelerated observer in $\mathrm{AdS}_{3}$ spacetime Jennings_2010 . It is demonstrated in Ref. Jennings_2010 that there exists a critical acceleration, $1/\ell$, in $\mathrm{AdS}_{3}$ spacetime, and only the observer with a constant super-critical acceleration $a$ (i.e., $a>1/\ell$) can register the quasi-thermal response with a temperature equal to $\sqrt{a^{2}-\ell^{-2}}/(2\pi)$. Besides, from the response functions $\gamma_{\pm}$ in (3) we can see that the Fermi-Dirac type distribution is displayed as noted in Ref. PhysRevD.49.1929 .
@@ 3.1 Charging dynamics
Let us turn our attention to analytically evaluating the average energy in some special cases. In the infinity time limit, i.e., $\tau\rightarrow\infty$, for both the decoherence and pure dephasing coupling cases the average energies in (3.1), (34), (3.1), (36) approaches to
@@
We note that in the infinity evolution time limit the average energies for the QB and charger are independent of the kinds of dissipation. Besides, in this case the average energies do not depend on the charging time, and is quite different from the case without dissipation shown in (30), which is alway periodic with the charging time.
@@
Figure 1: (a) The Hawking temperature shown in Eq. (29) as a function of the relative distance $r/r_{h}$ of the QB; (b) The average energy stored in the QB $\langle\,E_{\text{QB}}(\infty)\rangle$ shown in Eq. (3.1) (in units $\Delta\ell$) as a function of the relative distance $r/r_{h}$; (c) The average energy stored in the QB $\langle\,E_{\text{QB}}(\infty)\rangle$ shown in Eq. (3.1) (in units $\Delta\ell$) as a function of the charging amplitude $A\ell$.
@@
In Fig. 1 we show the Hawking temperature (29) and the average energy (3.1) shored in the QB in the infinity charging time. The temperature decreases monotonously as the increase of the distance of the QB away from the BTZ black hole horizon $r/r_{h}$. In the infinity charging time limit, the average energy shored in the QB approaches to $\lim_{r/r_{h}\rightarrow 1}\langle\,E_{\text{QB}}(\infty)\rangle=\frac{\Delta}{2}$ when the QB is located near the horizon, which does not depend on the charging amplitude $A$. This average energy decreases sharply with the increase of the relative distance $r/r_{h}$ when near the horizon. While far away from the horizon it approaches to a constant $\lim_{r/r_{h}\rightarrow\infty}\langle\,E_{\text{QB}}(\infty)\rangle=\frac{\Delta}{2}(1-\frac{\Delta}{\Omega})$ which depends on the charging amplitude. We can find that when QB far away from the horizon the stronger the charging amplitude is, the more the average energy shored in the QB is. It means that, in the presence of dissipation, larger driving amplitude $A$ leads to better charging of the QB. We can also see this behavior in Fig. 1 (c). Furthermore, the further the QB locates away from the black hole horizon, the more strongly its average energy in the infinite time limit depends on the charging amplitude $A$. This is because that when far way from the black hole horizon, the Hawking temperature in (29) approaches to vanishment, the dissipation of the QB effectively coming from the vacuum fluctuations of the quantum field could be considered to be weak. In this case the charger thus would play a dominated role in the whole charging process.
@@
When $A\geq\Delta$, we can always find that the average-energy difference in this case is smaller than zero, $\delta E\leq 0$. Which means that the dissipation from vacuum field fluctuations in the BTZ spacetime might causes a worse charging performance compared with the closed QB case. Besides, if $T\gg\Omega$, we can find in Eq. (39) $\delta E\simeq\frac{\Delta}{2}(1-2A^{2}/\Omega^{2})=\Delta(\Delta^{2}-A^{2})/2\Omega^{2}$. While for $T\ll\Omega$, we can find $\delta E\simeq\frac{\Delta}{2}(1-\Delta/\Omega-2A^{2}/\Omega^{2})$. In these two cases, whether the charging performance in the presence of dissipation is better or worse than the closed QB charging case depends on the QB energy-level spacing $\Delta$ and the charging amplitude $A$. In Fig. 2 we plot the average-energy difference shown in (39) as a function of the QB energy-level spacing $\Delta\ell$ and the charging amplitude $A\ell$. We can find that the location of the QB (or equivalently the Hawking temperature), $r$ (or $T$), affects the charging performance very much. When the QB is relatively far away from the horizon $r_{h}$, the average energy stored in QB in the presence of dissipation would never be more than that for the closed QB case. Thus when the Hawking temperature is relatively smaller, the dissipation from vacuum field fluctuations in the BTZ spacetime causes harmful influence on the charging performance. However, we can also find that when the QB approaches to the horizon, where the Hawking temperature is higher, the average energy shored in the QB in the presence of dissipation might be more than that for the closed QB case under the condition, $T\geq\Omega/\ln\big{[}\frac{\Delta\Omega+\Delta^{2}-A^{2}}{\Delta\Omega+A^{2}-\Delta^{2}}\big{]}$ with $\Delta>A$. It means in this case the dissipation from vacuum field fluctuations in the BTZ spacetime might enhance the charging performance. In other words, one can extract the energy from the vacuum fluctuations in curved space through dissipation in the charging process.

### v4

@@ Abstract
We consider how charging performances of a quantum battery, modeled as a two-level system, are influenced by the presence of vacuum fluctuations of a quantum field satisfying the Dirichlet, transparent, and Neumann boundary conditions in the BTZ spacetime. The quantum battery is subjected to an external static driving which works as a charger. Meanwhile, the quantum field is assumed to be coupled to both longitudinal and transverse spin components of the quantum battery including decoherence and pure dephasing mechanisms. Charging and discharging dynamics of the quantum battery are derived by extending the previous open quantum system approach in the relativistic framework to this more general scenario including both the driving and multiple coupling. Analytic expressions for the time evolution of the energy stored are presented. We find that when the driving amplitude is stronger/weaker than the energy-level spacing of the quantum battery the pure dephasing dissipative coupling results in better/worse charging performances than the decoherence dissipative coupling case. We also find that higher local Hawking temperature helps to improve the charging performance under certain conditions compared with the closed quantum battery case, implying the feasibility of energy extraction from vacuum fluctuations in curved spacetime via dissipation in charging protocol. Different boundary conditions for quantum field may lead to different charging performance. Furthermore, we also address the charging stability by monitoring the energy behaviour after the charging protocol has been switched off. Our study presents a general framework to investigate relaxation effects in curved spacetime, and reveals how spacetime properties and field boundary condition affect the charging process, which in turn may shed light on the exploration of the spacetime properties and thermodynamics via the charging protocol.
@@ 1 Introduction
The notion of QB was first introduced in the seminal work PhysRevE.87.042123 by Alicki and Fannes, showing entangling unitary controls extract in general more work than independent ones. Afterward, the development of different figures of merit, such as the charging time, the associated power, and the amount of extractable work, has been a major focus on the QB physics Campaioli:2023ndh ; Binder_2015 ; PhysRevLett.122.047702 ; PhysRevLett.111.240401 ; PhysRevLett.111.050601 ; PhysRevLett.131.030402 ; PhysRevLett.129.130602 ; Fischer . To improve the charging performance, a lot of methods have been developed to assist the charging protocol, including collective quantum resources PhysRevLett.120.117702 ; PhysRevLett.118.150601 ; PhysRevLett.128.140501 , ordered and disordered interactions PhysRevA.101.032115 , external environment engineering PhysRevA.107.032203 ; PhysRevE.104.064143 ; PhysRevA.109.022226 ; PhysRevA.102.052223 , collisional model PhysRevLett.127.100601 ; PhysRevResearch.5.013155 , Sachdev-Ye-Kitaev model PhysRevLett.125.236402 ; SYK and so on; see Ref. Campaioli:2023ndh for an extensive review and a comprehensive list of references. Beyond the theoretical analysis, the experimental implementation of various kinds of QBs and different charging protocols has also been reported recently Campaioli:2023ndh ; PhysRevB.108.L180301 ; PhysRevLett.131.240401 ; Hu_2022 ; PhysRevA.106.042601 ; batteries8050043 . Furthermore, since any systems are unavoidably coupled to the environment, at least subjected to vacuum fluctuations at the quantum scale, relaxation and dephasing would play a key role in the system dynamics. In this sense, the QB’s figures of merit, including the maximum average energy stored during the charging protocol, the charging power, the charging stability and so on, inevitably undergoes the relaxation and dephasing, which may even dominate the QB charging performance. It is thus of great importance to learn the charging performance in the presence of inevitable external environmental action Campaioli:2023ndh ; PhysRevB.99.035421 ; PhysRevA.102.060201 ; PhysRevA.103.033715 ; Carrega_2020 ; PhysRevLett.122.210601 ; PhysRevE.104.044116 ; PhysRevA.104.032207 .
@@
The outline of this paper is as follows. In section 2 we present the model of a single cell QB coupled to a vacuum fluctuations of quantum field. The charging performance indicators, such as the average energy stored in the QB, are calculated here. In section 3 we consider the dynamics of the average energy stored when the open QB is located in the BTZ spacetime, and explore how the driving amplitude, BTZ spacetime properties, the boundary conditions on quantum field, and different dissipative couplings affect the charging performance, including both the dynamics of the average energy stored and the charging stability. In section 4, we study the quantum battery physics in AdS spacetime and compare the results with that of the BTZ spacetime case. Finally, discussions and summary of the main results are present in section 5.
@@ 2.1 Open quantum battery model
where $\mu$ is a small coupling constant. $\theta$ here is to control the weight of the QB-field coupling that along the longitudinal (z) and transverse (x) directions. This interaction (4) captures both decoherence $(\theta=0)$ and pure dephasing $(\theta=\pi/2)$ processes. Note that in additional to the linear coupling between matter and field (4) which is commonly used in study of Unruh-DeWitt detector when learning the quantum field theory in curved spacetime, actually other various coupling forms, e.g., the derivative coupling K-J-Hinton_1983 ; P-G-Grove_1986 ; Juarez-Aubry_2014 ; Louko ; Jorma ; PhysRevD.98.065006 ; Erickson ; PhysRevD.104.025001 ; Mann , and quadratic coupling K-J-Hinton_1984 ; Norikazu-Suzuki_1997 ; doi:10.1142/S0217732302007545 ; PhysRevD.93.024019 ; PhysRevD.96.085012 ; PhysRevD.103.125021 , have also absorbed much attention. Besides, weak coupling is common for matter-field interaction in nature, while its strong and even ultrastrong coupling regime have been demonstrated in a variety of settings recently RevModPhys.90.021002 ; RevModPhys.91.025005 ; Kockum ; RevModPhys.92.025003 .
@@
Rotating $\tilde{\rho}_{\text{BQ}}(\tau)$ in Eq. (12) with state parameters (2.1), we can obtain the evolving state of the QB, $\rho_{\text{QB}}(\tau)=R^{\dagger}(\Theta)\tilde{\rho}_{\text{QB}}(\tau)R(\Theta)$. Note that the unavoidable coupling with an environment is responsible for dissipation, leading to relaxation and dephasing of the QB during the charging protocol, which are respectively characterized by the incoherent relaxation rate $\Gamma^{\text{r}}_{\theta}=\gamma_{+}+\gamma_{-}$ and the dephasing rate $\Gamma_{\theta}=\frac{1}{2}(\gamma_{+}+\gamma_{-}+4\gamma_{z})$ shown in Eq. (2.1). During the relaxation process, the population of the quantum state of the QB (the diagonal terms in the density matrix of quantum state) might vary with respect to the evolution time. However, the dephasing process would not change the population of the quantum state, but lead to the degradation of coherence (reducing the off-diagonal terms in the density matrix of quantum state). As a result of that, the environment-induced dissipation can inevitably affect the dynamics of the QB, thus the charging performance in terms of the figures of merit of the QB.
@@ 2.2 Charging performance indicators
Recall that for $\tau>0$, since we have assumed that the driving is static, we have $\dot{H}=0$, with an energy balance of the form
@@
Note that $r_{3}(\tau)$ and $r_{1}(\tau)$ are respectively related to the diagonal and off-diagonal terms of the density matrix of the QB’s state, whose dynamics are determined by the incoherent relaxation rate and the dephasing rate respectively. Besides, both the incoherent relaxation rate $\Gamma^{\text{r}}_{\theta}$ and the dephasing rate $\Gamma_{\theta}$ depend on the quantum field correlation function shown in (2.1), thus $r_{3}(\tau)$ and $r_{1}(\tau)$ depend on the spacetime properties due to the scattering off quantum field. Therefore, one can expect that the charging dynamics depends on the spacetime properties, and in turn one can also learn the spacetime properties from the charging physics.
@@ 3 Quantum battery in BTZ spacetime
To obtain the charging dynamics of the QB in the BTZ spacetime, we first need to know the Wightman function of the conformal scalar field which may capture the character of the quantum field vacuum fluctuations and plays an important role in the dynamics of the QB through the QB-field interaction. In this paper, we assume that the scalar field is in the Hartle-Hawking vacuum in the BTZ spacetime. It is interesting to note that the BTZ spacetime can be obtained by a topological identification of $\mathrm{AdS}_{3}$ spacetime. Therefore, using the method of images, the corresponding Wightman function in the BTZ spacetime for the conformal massless scalar field in the Hartle-Hawking vacuum can be obtained in terms of the corresponding Wightman function in $\mathrm{AdS}_{3}$ spacetime  [^3] Go beyond the standard quantization of quantum field, an alternative quantization in AdS/CFT has been introduced in Ref. AdsCFT . PhysRevD.49.1929 ,
@@
with $\Delta\phi=\phi-\phi^{\prime}$, and $\Delta\,t=t-t^{\prime}$. The parameter $\zeta\in\{1,0,-1\}$ respectively specifies the Dirichlet ($\zeta=1$), transparent ($\zeta=0$), and Neumann ($\zeta=-1$) boundary conditions satisfied by the field at spatial infinity PhysRevD.18.3565 . Actually, the Wightman function in the BTZ spacetime is derived in terms of the corresponding Wightman function in $\mathrm{AdS}_{3}$ spacetime as shown in (22). This is based on the feature of the BTZ solution that the solution arises from identifying points in $\mathrm{AdS}_{3}$ space, using the orbits of a spacelike Killing vector field PhysRevD.49.1929 . However, since $\mathrm{AdS}_{3}$ is not a globally hyperbolic spacetime, it means that, in order to have well defined quantum field theory, appropriate boundary conditions must be applied to the field at spatial infinity. Or information can escape or leak in through this time-like surface in a finite coordinate time, spoiling the composition law property of the field propagator. This problem carries over to the black hole solution PhysRevD.49.1929 . Usually, the simplest boundary conditions are either Dirichlet (where the field vanishes on the boundary) or Neumann (where the normal derivative of the field vanishes on the boundary) PhysRevD.49.1929 ; PhysRevD.18.3565 , which means that information is reflected and not lost at the spatial infinity. It is also possible to define a quantization scheme on $\mathrm{AdS}_{3}$ space without using boundary conditions, which is referred to as “transparent” boundary conditions” in Ref. PhysRevD.18.3565 . However, in such case an “effective Cauchy surface” for $\mathrm{AdS}$ is needed to be constructed to recirculate the energy, angular momentum, etc., lost to timelike infinity, resulting in a well-defined, if rather unusual, conservation law (see more details in Ref. PhysRevD.18.3565 ). Different boundary conditions mean different physical properties of spatial infinity, quantum field at which behaves quite differently due to spacetime scattering, leading to different physical phenomena, e.g., significant effect on the transition rate for Unruh-DeWitt particle detector PhysRevD.86.064031 and the Fischer information FIBH . Therefore, it is of interest to study how different boundary conditions affect charging performance of the QB, and in turn, how to use such effects to explore the properties of spacetime.
@@
Let us note that the local temperature of the BTZ spacetime (we call this temperature the local Hawking temperature) can be rewritten in the following form
@@
with $a_{\text{BTZ}}=r/(\ell\sqrt{r^{2}-r^{2}_{\text{h}}})$ representing the acceleration of the constant $r$ trajectory in the BTZ spacetime. Here $T$ is analogous to the temperature felt by an accelerated observer in $\mathrm{AdS}_{3}$ spacetime Jennings_2010 . It is demonstrated in Ref. Jennings_2010 that there exists a critical acceleration, $1/\ell$, in $\mathrm{AdS}_{3}$ spacetime, and only the observer with a constant super-critical acceleration $a$ (i.e., $a>1/\ell$) can register the quasi-thermal response with a temperature equal to $\sqrt{a^{2}-\ell^{-2}}/(2\pi)$. Besides, from the response functions $\gamma_{\pm}$ in (3) we can see that the Fermi-Dirac type distribution is displayed as noted in Ref. PhysRevD.49.1929 . The response functions in (3) also are the function of various parameters of the QB itself and the BTZ black hole listed in the Table 1.
$\mu$ the coupling constant between the QB and field
$\Delta$ the QB’s energy-level spacing
$A$ the charging amplitude characterizing charing strength of the charger
$T=\frac{r_{\text{h}}}{2\pi\ell\sqrt{r^{2}-r^{2}_{\text{h}}}}$ the local KMS temperature of BTZ black hole
$r_{\text{h}}=\sqrt{M}\ell$ the horizon radius with $M$ being the mass of BTZ black hole
$\Lambda=-1/\ell^{2}$ the negative cosmological constant
$r$ the distance of the observer (or QB) away from the BTZ black hole
$\zeta$ the boundary condition parameter
Table 1: Various parameters of both the QB and BTZ black hole. All of them will affect the dynamics of the QB and thus the charging performance.
Clearly, these parameters determine the dissipation and thus affect the dynamics of the QB. In what follows, how the different parameters affect charging performance of the QB will be studied.
@@ 3.1 Charging dynamics
Let us turn our attention to analytically evaluating the average energy in some special cases. In the infinite time limit, i.e., $\tau\rightarrow\infty$, for both the decoherence and pure dephasing coupling cases the average energies in (3.1), (34), (3.1), (36) approach to
@@
We note that in the infinite evolution time limit the average energies for the QB and charger are independent of the kinds of dissipation. Besides, in this case the average energies do not depend on the charging time, and is quite different from the case without dissipation shown in (30), which is alway periodic with the charging time.
@@
Figure 1: (a) The local KMS temperature of BTZ black hole shown in Eq. (29) as a function of the relative distance $r/r_{h}$ of the QB; (b) The average energy stored in the QB $\langle\,E_{\text{QB}}(\infty)\rangle$ shown in Eq. (3.1) (in units $\Delta\ell$) as a function of the relative distance $r/r_{h}$; (c) The average energy stored in the QB $\langle\,E_{\text{QB}}(\infty)\rangle$ shown in Eq. (3.1) (in units $\Delta\ell$) as a function of the charging amplitude $A\ell$.
@@
In Fig. 1 we show the local KMS temperature of BTZ black hole (29) and the average energy (3.1) stored in the QB in the infinite charging time. The temperature decreases monotonously as the increase of the distance of the QB away from the BTZ black hole horizon $r/r_{\text{h}}$. It means that the further the QB is located away from the black horizon, the lower the temperature that the QB feels is. In the open quantum system theory, the thermal bath with higher temperature will lead to stronger dissipation of the QB (see more details below). In the infinite charging time limit, the average energy shored in the QB approaches to $\lim_{r/r_{h}\rightarrow 1}\langle\,E_{\text{QB}}(\infty)\rangle=\frac{\Delta}{2}$ when the QB is located near the horizon, which does not depend on the charging amplitude $A$. This suggests that when the temperature of thermal bath (or local Hawking temperature) is extremely high the stationary average energy that is deposited in the QB through charging only depends on its energy-level spacing, while the other information of external environment, such as the charging amplitude and black hole background parameters, is not encoded in the QB in this scenario. Besides, this average energy decreases sharply with the increase of the relative distance $r/r_{\text{h}}$ when near the horizon. That is to say, in the extremely high-temperature thermal bath, even small temperature fluctuation would lead to strong dissipation and thus the drastic change of the average energy stored in the QB. While when the QB is far way from the horizon (the local Hawking temperature that the QB feels approaches to zero), the average energy approaches to a constant $\lim_{r/r_{\text{h}}\rightarrow\infty}\langle E_{\text{QB}}(\infty)\rangle=\frac{\Delta}{2}(1-\frac{\Delta}{\Omega})$. It depends on the charging amplitude and is different from the case near the horizon. We can find that when QB far away from the horizon the stronger the charging amplitude is, the more the average energy stored in the QB is. This means that, in the presence of dissipation from vacuum fluctuations of external environment, larger driving amplitude $A$ leads to batter charging of the QB. We can also see this behavior in Fig. 1 (c). Furthermore, the further the QB locates away from the black hole horizon, the more strongly its average energy in the infinite time limit depends on the charging amplitude $A$. This is because that when far way from the black hole horizon, the local Hawking temperature in (29) approaches to vanishment, the dissipation of the QB effectively coming from the vacuum fluctuations of the quantum field could be considered to be weak. In this case the charger thus would play a dominated role in the whole charging process.
@@
When $A\geq\Delta$, we can always find that the average-energy difference in this case is smaller than zero, $\delta E\leq 0$. Which means that the dissipation from vacuum field fluctuations in the BTZ spacetime might causes a worse charging performance compared with the closed QB case. Besides, if $T\gg\Omega$, we can find in Eq. (39) $\delta E\simeq\frac{\Delta}{2}(1-2A^{2}/\Omega^{2})=\Delta(\Delta^{2}-A^{2})/2\Omega^{2}$. While for $T\ll\Omega$, we can find $\delta E\simeq\frac{\Delta}{2}(1-\Delta/\Omega-2A^{2}/\Omega^{2})$. In these two cases, whether the charging performance in the presence of dissipation is better or worse than the closed QB charging case depends on the QB energy-level spacing $\Delta$ and the charging amplitude $A$. In Fig. 2 we plot the average-energy difference shown in (39) as a function of the QB energy-level spacing $\Delta\ell$ and the charging amplitude $A\ell$. We can find that the location of the QB (or equivalently the local Hawking temperature), $r$ (or $T$), affects the charging performance very much. When the QB is relatively far away from the horizon $r_{h}$, the average energy stored in QB in the presence of dissipation would never be more than that for the closed QB case. Thus when the local Hawking temperature is relatively smaller, the dissipation from vacuum field fluctuations in the BTZ spacetime causes harmful influence on the charging performance. However, we can also find that when the QB approaches to the horizon, where the local Hawking temperature is higher, the average energy shored in the QB in the presence of dissipation might be more than that for the closed QB case under the condition, $T\geq\Omega/\ln\big{[}\frac{\Delta\Omega+\Delta^{2}-A^{2}}{\Delta\Omega+A^{2}-\Delta^{2}}\big{]}$ with $\Delta>A$. It means in this case the dissipation from vacuum field fluctuations in the BTZ spacetime might enhance the charging performance. In other words, one can extract the energy from the vacuum fluctuations in curved space through dissipation in the charging process.